Government Expenditure: 8
Consumption Expenditure: 35
Disposable Income: 50
Net Taxes: 12
Imports: 7
Exports: 9

Data is in millions of US dollars.

a. Based on the data provided above, what is the value of National Income (GDP or Y)?
b. Based on the data provided above, what is the value of Private Savings?
c. Based on the data provided above, what is the value of Government Savings?
d. Based on the data provided above, what is the value of National Savings?
e. Based on the data provided above, what is the value of Net Exports?
f. Based on the data provided above, is this country a net lender or net borrower?
1
Expert's answer
2019-08-15T09:19:50-0400

1) Gross domestic product (GDP)=Government expenditure+Consumption expenditure+(export-import)

=8+35+9-7

=45 millions dollar

2)Private saving=Disposable income-Consumption

=50-35

=15 million dollars

3) Government savings=Taxes-Government expenditure

=12-8

=4 millons dollar

4) National savings= Private saving+Government savings

=15+4

=19 million dollars

5) Net exports= exports-imports

=9-7

=2 million dollars

6) As net export is positive means export is more than import. It shows that country is lender